# FDS: Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization

### v2

@@ Appendix D Detailed Results
Method Aug. Target Domains Art Clipart Product Real World Avg. Standard Methods ERM (baseline) [44] 61.3 $\pm$0.7 52.4 $\pm$0.3 75.8 $\pm$0.1 76.6 $\pm$0.3 66.5 $\pm$0.3 ERM (reproduced) 59.5 $\pm$2.1 51.3 $\pm$1.3 73.8 $\pm$0.8 73.8 $\pm$0.2 64.6 $\pm$1.1 IRM [60] 58.9 $\pm$2.3 52.2 $\pm$1.6 72.1 $\pm$2.9 74.0 $\pm$2.5 64.3 $\pm$2.2 GroupDRO [47] 60.4 $\pm$0.7 52.7 $\pm$1.0 75.0 $\pm$0.7 76.0 $\pm$0.7 66.0 $\pm$0.7 Mixup [61] 62.4 $\pm$0.8 54.8 $\pm$0.6 76.9 $\pm$0.3 78.3 $\pm$0.2 68.1 $\pm$0.3 CORAL [62] 65.3 $\pm$0.4 54.4 $\pm$0.5 76.5 $\pm$0.1 78.4 $\pm$0.5 68.7 $\pm$0.3 MMD [63] 60.4 $\pm$0.2 53.3 $\pm$0.3 74.3 $\pm$0.1 77.4 $\pm$0.6 66.3 $\pm$0.1 DANN [64] 59.9 $\pm$1.3 53.0 $\pm$0.3 73.6 $\pm$0.7 76.9 $\pm$0.5 65.9 $\pm$0.6 MLDG [15] 61.5 $\pm$0.9 53.2 $\pm$0.6 75.0 $\pm$1.2 77.5 $\pm$0.4 66.8 $\pm$0.7 VREx [74] 60.7 $\pm$0.9 53.0 $\pm$0.9 75.3 $\pm$0.1 76.6 $\pm$0.5 66.4 $\pm$0.6 ARM [75] 58.9 $\pm$0.8 51.0 $\pm$0.5 74.1 $\pm$0.1 75.2 $\pm$0.3 64.8 $\pm$0.4 SagNet [65] 63.4 $\pm$0.2 54.8 $\pm$0.4 75.8 $\pm$0.4 78.3 $\pm$0.3 68.1 $\pm$0.1 RSC [23] 60.7 $\pm$1.4 51.4 $\pm$0.3 74.8 $\pm$1.1 75.1 $\pm$1.3 65.5 $\pm$0.9 Mixstyle [20] 51.1 $\pm$0.3 53.2 $\pm$0.4 68.2 $\pm$0.7 69.2 $\pm$0.6 60.4 $\pm$0.3 mDSDI [66] 68.1 $\pm$0.3 52.1 $\pm$0.4 76.0 $\pm$0.2 80.4 $\pm$0.2 69.2 $\pm$0.4 SelfReg [41] 63.6 $\pm$1.4 53.1 $\pm$1.0 76.9 $\pm$0.4 78.1 $\pm$0.4 67.9 $\pm$0.7 Fishr [73] 62.4 $\pm$0.5 54.4 $\pm$0.4 76.2 $\pm$0.5 78.3 $\pm$0.1 67.8 $\pm$0.5 DCAug [59] 61.8 $\pm$0.6 55.4 $\pm$0.6 77.1 $\pm$0.3 78.9 $\pm$0.3 68.3 $\pm$0.4 CDGA [55] 60.1 $\pm$1.4 54.2 $\pm$0.5 78.2 $\pm$0.6 80.4 $\pm$0.1 68.2 $\pm$0.6 ERM + FDS (ours) 64.6 $\pm$0.2 57.7 $\pm$0.1 80.2 $\pm$0.5 82.0 $\pm$0.4 71.1 $\pm$0.1 WA Methods SWAD (baseline) [24] 66.1 $\pm$0.4 57.7 $\pm$0.4 78.4 $\pm$0.1 80.2 $\pm$0.2 70.6 $\pm$0.2 SWAD (reproduced) 65.9 $\pm$0.9 56.8 $\pm$0.4 78.8 $\pm$0.3 80.0 $\pm$0.2 70.3 $\pm$0.4 SelfReg SWA [41] 64.9 $\pm$0.8 55.4 $\pm$0.6 78.4 $\pm$0.2 78.8 $\pm$0.1 69.4 $\pm$0.2 DNA [67] 67.7 $\pm$0.2 57.7 $\pm$0.3 78.9 $\pm$0.2 80.5 $\pm$0.2 71.2 $\pm$0.1 DiWA [68] 67.3 $\pm$0.2 57.9 $\pm$0.2 79.0 $\pm$0.2 79.9 $\pm$0.1 71.0 $\pm$0.1 TeachDCAug [59] 66.2 $\pm$0.2 57.0 $\pm$0.3 78.3 $\pm$0.1 80.1 $\pm$0.0 70.4 $\pm$0.2 SWAD + FDS (ours) 67.3 $\pm$0.8 60.5 $\pm$0.5 82.6 $\pm$0.1 83.6 $\pm$0.3 73.5 $\pm$0.4

### v3

@@ Appendix D Detailed Results
Method Aug. Target Domains Art Clipart Product Real World Avg. Standard Methods ERM (baseline) [44] 61.3 $\pm$0.7 52.4 $\pm$0.3 75.8 $\pm$0.1 76.6 $\pm$0.3 66.5 $\pm$0.3 ERM (reproduced) 59.5 $\pm$2.1 51.3 $\pm$1.3 73.8 $\pm$0.8 73.8 $\pm$0.2 64.6 $\pm$1.1 IRM [61] 58.9 $\pm$2.3 52.2 $\pm$1.6 72.1 $\pm$2.9 74.0 $\pm$2.5 64.3 $\pm$2.2 GroupDRO [47] 60.4 $\pm$0.7 52.7 $\pm$1.0 75.0 $\pm$0.7 76.0 $\pm$0.7 66.0 $\pm$0.7 Mixup [62] 62.4 $\pm$0.8 54.8 $\pm$0.6 76.9 $\pm$0.3 78.3 $\pm$0.2 68.1 $\pm$0.3 CORAL [63] 65.3 $\pm$0.4 54.4 $\pm$0.5 76.5 $\pm$0.1 78.4 $\pm$0.5 68.7 $\pm$0.3 MMD [64] 60.4 $\pm$0.2 53.3 $\pm$0.3 74.3 $\pm$0.1 77.4 $\pm$0.6 66.3 $\pm$0.1 DANN [65] 59.9 $\pm$1.3 53.0 $\pm$0.3 73.6 $\pm$0.7 76.9 $\pm$0.5 65.9 $\pm$0.6 MLDG [14] 61.5 $\pm$0.9 53.2 $\pm$0.6 75.0 $\pm$1.2 77.5 $\pm$0.4 66.8 $\pm$0.7 VREx [75] 60.7 $\pm$0.9 53.0 $\pm$0.9 75.3 $\pm$0.1 76.6 $\pm$0.5 66.4 $\pm$0.6 ARM [76] 58.9 $\pm$0.8 51.0 $\pm$0.5 74.1 $\pm$0.1 75.2 $\pm$0.3 64.8 $\pm$0.4 SagNet [66] 63.4 $\pm$0.2 54.8 $\pm$0.4 75.8 $\pm$0.4 78.3 $\pm$0.3 68.1 $\pm$0.1 RSC [22] 60.7 $\pm$1.4 51.4 $\pm$0.3 74.8 $\pm$1.1 75.1 $\pm$1.3 65.5 $\pm$0.9 Mixstyle [19] 51.1 $\pm$0.3 53.2 $\pm$0.4 68.2 $\pm$0.7 69.2 $\pm$0.6 60.4 $\pm$0.3 mDSDI [67] 68.1 $\pm$0.3 52.1 $\pm$0.4 76.0 $\pm$0.2 80.4 $\pm$0.2 69.2 $\pm$0.4 SelfReg [41] 63.6 $\pm$1.4 53.1 $\pm$1.0 76.9 $\pm$0.4 78.1 $\pm$0.4 67.9 $\pm$0.7 Fishr [74] 62.4 $\pm$0.5 54.4 $\pm$0.4 76.2 $\pm$0.5 78.3 $\pm$0.1 67.8 $\pm$0.5 DCAug [60] 61.8 $\pm$0.6 55.4 $\pm$0.6 77.1 $\pm$0.3 78.9 $\pm$0.3 68.3 $\pm$0.4 DomainDiff [54] 57.6 $\pm$0.4 49.2 $\pm$0.6 73.0 $\pm$0.6 75.2 $\pm$0.9 63.7 $\pm$0.6 CDGA [56] 60.1 $\pm$1.4 54.2 $\pm$0.5 78.2 $\pm$0.6 80.4 $\pm$0.1 68.2 $\pm$0.6 ERM + FDS (ours) 64.6 $\pm$0.2 57.7 $\pm$0.1 80.2 $\pm$0.5 82.0 $\pm$0.4 71.1 $\pm$0.1 WA Methods SWAD (baseline) [23] 66.1 $\pm$0.4 57.7 $\pm$0.4 78.4 $\pm$0.1 80.2 $\pm$0.2 70.6 $\pm$0.2 SWAD (reproduced) 65.9 $\pm$0.9 56.8 $\pm$0.4 78.8 $\pm$0.3 80.0 $\pm$0.2 70.3 $\pm$0.4 SelfReg SWA [41] 64.9 $\pm$0.8 55.4 $\pm$0.6 78.4 $\pm$0.2 78.8 $\pm$0.1 69.4 $\pm$0.2 DNA [68] 67.7 $\pm$0.2 57.7 $\pm$0.3 78.9 $\pm$0.2 80.5 $\pm$0.2 71.2 $\pm$0.1 DiWA [69] 67.3 $\pm$0.2 57.9 $\pm$0.2 79.0 $\pm$0.2 79.9 $\pm$0.1 71.0 $\pm$0.1 TeachDCAug [60] 66.2 $\pm$0.2 57.0 $\pm$0.3 78.3 $\pm$0.1 80.1 $\pm$0.0 70.4 $\pm$0.2 SWAD + FDS (ours) 67.3 $\pm$0.8 60.5 $\pm$0.5 82.6 $\pm$0.1 83.6 $\pm$0.3 73.5 $\pm$0.4
@@
Figure 13: t-SNE visualization of all classes from the PACS dataset, showing the distribution of original source domains (Art, Photo, Sketch) and FDS ones.
Figure 14: Visual comparison of original "Art" and "Photo" samples from PACS with synthetic images generated by FDS (Art + Photo). The middle section illustrates how FDS combines visual elements from both domains, producing diverse, domain-bridging images.
Figure 15: Visual comparison of original "Sketch" and "Art" samples from PACS with synthetic images generated by FDS (Sketch + Art). The generated images in the middle section showcase a blend of artistic textures and sketched outlines.
Figure 16: Visual comparison of original "Photo" and "Sketch" samples from PACS with synthetic images generated by FDS (Photo + Sketch). The middle section demonstrates how FDS integrates the photorealistic details of the "Photo" domain with the elements of the "Sketch" domain.